The coordinates of the point (3,-5) in the new system when the origin is shifted to (-23) are

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
lets take new origin to be
(h, k) = (-2, 3) be the given point
 
 so let the new co-ordinates (X, Y) .
x = X + h and y = Y + k (shifting of origin formula)
 
i.e. 3 = X + (-2) and -5 = Y + 3
 
This gives, X = 5 and Y = -8.
 
Thus the new co-ordinates are (5,-8)  - Ans.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)